Solve the following for b
b-3=-2(x+9)

Answer by ewatrrr(24785)


Hi
Solving the following for b:
b-3=-2(x+9)
b = -2(x+9) + 3
b = -2x - 18 + 3
b = -15-2x or b = -(15+2x)